OpenStudy (anonymous):

I give medals 2x + y = 3 x - 2y = -1 If equation one is multiplied by 2 and then the equations are added, the result is _____. 3x = 2 3x = 5 5x = 5

OpenStudy (anonymous):

Just use system of equations elimination or substituion

OpenStudy (math&ing001):

2 * [(2x + y) = 3] So that gives you : 4x +2y = 6 + x - 2y = -1 Do the math

OpenStudy (anonymous):

um so the answer is b right? or am i wrong

OpenStudy (math&ing001):

Nope it's C. 4x+x=5x

OpenStudy (anonymous):

damn. okay thanks